OpenStudy (anonymous):

Find the slope of the straight line that passes through the given points or state that the slope is undefined. Then indicate if the line through the points rises (left to right), falls (left to right), is horizontal, or is vertical 1a. (2, -5) and (-1, -1) 1b. (7, 1) and (7, 2) 1c. (1, 7) and (-2, -2) 1d. (-1, -5) and (-4, -5)

OpenStudy (mimi_x3):

lols , use the gradient formula y^2 - y^2 / x^2 - x^1

OpenStudy (anonymous):

in each case compute \[m=\frac{y_2-y_1}{x_2-x_1}\] so for example the first one would be \[,=\frac{-1+5}{-1-2}=-\frac{4}{3}\]

OpenStudy (anonymous):

don't know where i got that from. second one x values are the same, so slope does not exist

OpenStudy (anonymous):

third one is the one i wrote and deleted: \[m=\frac{-2-7}{-2-1}=\frac{-9}{-3}=3\]

OpenStudy (anonymous):

4th one y values are the same so slope is 0

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw